# Physically Feasible Semantic Segmentation

### v3

@@ 4.4 Ablation Study
Table 5: Feasible and infeasible inclusion pair statistics for different datasets. The statistics are computed on the training set of each dataset. Infeasible pair counts exclude the non-cooccurring pairs.
@@
4.5 Empirical Analysis
The feasible inclusion pairs are computed offline individually from the training set of each dataset using Algorithm 1. The complete count of the feasible and infeasible pairs of inclusion are given in Table 5. Every dataset is divided into train/val/test set. Algorithm 1 is applied to the train set of every dataset even when we jointly re-trained the models on $train+val$ sets.
5 Conclusion and Future Work
In this paper, rather than building a new semantic segmentation architecture, we have attempted to enhance any existing state-of-the-art network by enforcing a physical, data-driven prior into the network. Our method, PhyFea, constitutes a light-weight, plug-and-play module for semantic segmentation, which can be coupled with any network architecture to improve accuracy and conformity to our exemplary inclusion prior. We thus believe it is a first yet firm step towards physically-informed semantic perception. Overall, we are successful in attaining our goal of physical feasibility in segmentation both in the qualitative and quantitative sense.
Supplementary Material
6 Additional Qualitative Comparisons
Figure 8: Additional qualitative comparison on ADE20K. From left to right: input image, ground-truth semantic labels, and predictions of OCRNet [45] and PhyFea. Best viewed on a screen and zoomed in.
ADE20K.  In Fig. 8 showing additional qualitative results on the ADE20K val set, we observe that in the prediction of the basic OCRNet network on the first example building includes sky, which is rectified in the PhyFea-upgraded version of the network. In the second example, we observe that a segment of bed is infeasibly included in a cupboard segment, which is corrected by our PhyFea model.
Figure 9: Additional qualitative comparison on ACDC. From left to right: input image, ground-truth semantic labels, and predictions of SegFormer [41] and PhyFea. Best viewed on a screen and zoomed in.
ACDC.  In the first example of Fig. 9, we observe that a segment of sky is completely included in a building, which is a physical anomaly and is solved by PhyFea.
Figure 10: Additional qualitative comparison on Cityscapes. From left to right: input image, ground-truth semantic labels, and predictions of Mask2Former [6] and PhyFea. Best viewed on a screen and zoomed in.
Cityscapes.  In Fig. 10, we note that a rider segment corresponding to the rider’s arm is infeasibly included in a bus segment in the prediction of Mask2Former [6], whereas PhyFea manages to correctly connect this part of the arm with the rest of the rider’s body. In the second example, a sky segment is infeasibly included in a building segment in the Mask2Former prediction, while PhyFea avoids this error and better segments distant people which appear very small in the image.
7 Analysis of Prediction Feasibility
We define a metric named mean infeasibility normalized frequency (mINF) on an evaluation set using:
$\centering\text{mINF}=\frac{1}{|\mathcal{C}^{\prime}|}\sum_{(c_{i},c_{j})\in\mathcal{C}^{\prime}}\frac{f_{\text{inclusion}}(c_{i},c_{j})}{f_{\text{co-occur}}(c_{i},c_{j})}\@add@centering$ (6)
We compute $f_{\text{co-occur}}(c_{i},c_{j})$, i.e., the frequency of co-occurrence of the classes in the infeasible class pair $(c_{i},c_{j})$ in the set of predicted labelings on the evaluation set. $f_{\text{inclusion}}(c_{i},c_{j})$ is the frequency of the occurrence of infeasible inclusions for class pair $(c_{i},c_{j})$ in the set of predicted labelings on the evaluation set. An mINF of 0 signifies the complete absence of infeasible inclusions in the predicted labelings. We report mINF scores in Tab. 6 for SegFormer and Mask2Former models on Cityscapes and ACDC. These results show that PhyFea can almost entirely solve the anomaly of infeasible Inclusions. It achieved an mINF of almost 0.
Table 6: Comparison of prediction feasibility with state-of-the-art models on Cityscapes and ACDC. Results are reported with our mINF metric (%, lower is better) on the Cityscapes test set and the ACDC test set.
@@ 7 Analysis of Prediction Feasibility
PhyFea w/ SegFormer (ours) 0.1 0.2
@@
PhyFea w/ Mask2Former (ours) 0.1 0.1

### v4

@@ 4.4 Evaluation of Prediction Feasibility
Table 5: Feasible and infeasible inclusion pair statistics for different datasets. The statistics are computed on the training set of each dataset. Infeasible pair counts exclude pairs that do not co-occur in any image of the dataset.
@@
Table 6: Comparison of prediction feasibility with state-of-the-art models on Cityscapes and ACDC. Results are reported with our mINF metric (%, lower is better) on the Cityscapes and ACDC test sets.
@@
InSeIn w/ SegFormer (ours) 0.1 0.2
@@
InSeIn w/ Mask2Former (ours) 0.1 0.1
OneFormer [25] 3.0 7.1
InSeIn w/ OneFormer (ours) 0.1 0.1
Table 7: False response error (%, lower is better) on the Cityscapes and ACDC validation sets. We have set boundary regions to 15 pixels.
Method \ Dataset Cityscapes ACDC
SegFormer [44] 23.7 37.6
InSeIn w/ SegFormer (ours) 22.2 35.9
Mask2Former [9] 22.0 29.3
InSeIn w/ Mask2Former (ours) 20.8 28.0
OneFormer [25] 21.7 28.3
InSeIn w/ OneFormer (ours) 19.9 26.8
We compute $f_{\text{co-occur}}(c_{i},c_{j})$, i.e., the frequency of co-occurrence of the classes in the infeasible class pair $(c_{i},c_{j})$ in the set of predicted labelings on the evaluation set. $f_{\text{inclusion}}(c_{i},c_{j})$ is the frequency of the occurrence of infeasible inclusions for class pair $(c_{i},c_{j})$ in the set of predicted labelings on the evaluation set. An mINF of 0 signifies the complete absence of infeasible inclusions in the predicted labelings. We report mINF scores in Tab. 6 for SegFormer, Mask2Former, and OneFormer models on Cityscapes and ACDC. These results show that InSeIn can almost fully adhere to the inclusion constraint imposed by our inclusion loss.
We have also computed another pixel-level metric, namely the false response error [3, 51], with results for three representative network architectures shown in Tab. 7. Lower values imply that the prediction feature maps contain less erroneous pixels. We observe a significant improvement with InSeIn over the baseline predictions across all three examined architectures.
We additionally conduct an “oracle” experiment, by manually filtering with a conceptual process the original set of infeasible class pairs extracted from Cityscapes, which contains 180 pairs. After this filtering, the number of infeasible pairs becomes 120. The mIoU on the Cityscapes test set of the InSeIn model trained with this updated set of infeasible pairs based on a Mask2Former architecture is 83.9%, which is on a par with the original score presented in Tab. 1. This confirms that our data-driven constraint extraction already provides a meaningful set of constraints for moderate-size datasets such as Cityscapes, with minimal to no impact from potential spurious infeasibility constraints that are included to our loss due to the rare occurrence of the respective configurations in practice.
4.5 Ablation Study
Tab. 8 ablates InSeIn on Cityscapes using Mask2Former as the baseline. We report validation set performance. We adjust the padding appropriately to preserve spatial dimensions across dilation iterations. We observe that by increasing the kernel size with a decrease in the number of iterations for max-pooling, performance tends to decrease. The reason behind this decrease is that a bigger structuring element, i.e., the max-pooling kernel, enables the propagation of the border pixels to the interior portion of the feature map, which should otherwise be disconnected, and as a result, the infeasible included segment is not opened correctly. Thus, inclusion loss will assume too high a value, penalizing feasible segments.
5 Conclusion and Future Work
Table 8: Ablation study of InSeIn with Mask2Former-SwinB on the Cityscapes validation set. Number of iterations used for max-pooling kernels $5\times 5$ and $7\times 7$ is 60k.
Kernel-size Stride Padding $T$ mIoU
$3\times 3$ 1 1 128 85.3
$5\times 5$ 1 2 60 81.8
$7\times 7$ 1 3 3 79.8
In this paper, rather than building a new semantic segmentation architecture, we have attempted to enhance any existing state-of-the-art network by enforcing a data-driven high-level constraint into the network. Our method, InSeIn, constitutes a light-weight, plug-and-play module for semantic segmentation, which can be coupled with any network architecture to improve accuracy and conformity to our exemplary inclusion prior. Overall, InSeIn successfully attains our goal of structural feasibility in segmentation, both in the qualitative and quantitative sense.